# A variational formulation of stochastic thermodynamics. Part I: Finite-dimensional systems

### v1

@@ III.3 Mechanical isolated system
where $\mathbf{f_{-}}=-\underline{\underline{\bm{\gamma}}}(\mathbf{x},\mathbf{p},s)\mathbf{p}/m$ is a dissipative force with state-dependent damping coefficient, $\underline{\underline{\mathbf{g}}}(\mathbf{x},\mathbf{p},s)$ is the noise strength and $\bm{\zeta}(t)$ is gaussian white noise verifying the relations given by Eq. (2). $\Gamma$ denotes the (generalized) thermal displacement, i.e. $\dot{\Gamma}=T$, an equality that follows from the variational principle.
@@
with $\dot{s}=\dot{\Sigma}$, equality that holds in any closed system. See Appendix B for the detailed derivation. Taking the total time derivative of $E=\frac{1}{2}m|\mathbf{v}|^{2}+U(\mathbf{x},s)$ we find the first law
@@
$j_{d}=$ $\left(-\gamma\frac{p}{m}-g\partial_{p}g+\frac{p}{m}g\partial_{s}(g/T)\right)\mathcal{P}$ (45a)
$-g^{2}\partial_{p}\mathcal{P}+\frac{pg^{2}}{mT}\partial_{s}\mathcal{P},$
@@
the system and thermodynamic average EPRs yield:
$\dot{\mathcal{S}}=\int d\Omega\,\left\{\frac{{j}_{d}^{2}}{\mathcal{P}g^{2}}\right$.
@@
$\langle\!\langle\dot{s}\rangle\!\rangle=-\int d\Omega\,\frac{p}{mT}{j}_{d}.$ (47)
Enforcing the second law we then derive the FDR:
$\dot{S}_{tot}=\dot{\mathcal{S}}+\langle\!\langle\dot{s}\rangle\!\rangle=\int d\Omega\,\frac{{j}_{d}^{2}}{\mathcal{P}D}\geq 0\iff$ (48)
$-\gamma\frac{p}{m}=-\frac{D}{T}\frac{p}{m}+g\partial_{p}g-g\partial_{s}(g/T)\frac{p}{m},$ (49)
where we recall $D=g^{2}$. The total EPR has the same form as that obtained from the standard ST approach in one dimension, see Eq. (23), but it is instead defined over the thermodynamic phase space $\Omega$. The corresponding multidimensional version of Eq. (49), obtained by an analogous computation, is
@@
Remarkably, the thermodynamic EP satisfies
@@
which implies that $\dot{S}_{tot}$ is a real time derivative, as we can write
$\dot{S}_{tot}=\frac{d}{dt}(\mathcal{S}+\langle s\rangle),$ (52)
in contrast with the standard ST setting, i.e. the results of Sec. II.
Proceeding to analyze the FDR in Eq. (49), we observe that it connects the dissipative forces with the noise amplitude, and also includes contributions from noise-induced drift terms. In the special case of constant temperature and constant damping coefficient, we recover the standard Einstein relation discussed in Sec. II.3, namely $T\gamma=D$.
When $\gamma,T,g$ are independent of the entropy $s$, our result reproduces the FDR derived in Ref. [52], where this relation was derived by setting the steady-state distribution to be the Boltzmann distribution. This reference also illustrates how momentum-dependent multiplicative noise can arise, for instance, in systems with Coulomb friction, which requires an extension of the FDR to ensure thermodynamic consistency. Additionally, in this scenario—or more generally, for $\partial_{p}{\rm f}_{+}=0$—where the observable DoFs decouple from the entropy balance in (LABEL:eq:EoM_mech_closed), $\dot{s}$ recovers the same expression as $\dot{s}_{m}$ in Eq. (18) under the FDR constraint, providing it with a well-defined thermodynamic interpretation (Clausius relation).
Beyond recovering these known results, this framework introduces a novel approach for coupling mechanical and thermal DoFs [24]. This allows for taking into account finite baths and/or coupled system-bath dynamics, thus going beyond the weak coupling limit. The third contribution in Eq. (49), the noise-induced drift due to $s$ dependence, is essential to guarantee thermodynamic consistency in the presence of such nonlinear system-bath coupling.
@@
which implies $\partial_{t}\mathcal{P}=-\nabla_{\Omega}\cdot\mathbf{j}=0$. The distribution (53) is general and holds for coupled system-bath dynamics as well. Moreover, it favors the states that maximize the entropy $s$, consistent with the maximum entropy principle in statistical mechanics, as well as being of the same form as the equilibrium distribution in fluctuation theory [49, 53, 54, 55], with an additional prefactor accounting for energy conservation. The correspondence between this equilibrium state and fluctuation theory arises since both assume gaussian fluctuations.
@@
This corresponds precisely to the Maxwell-Boltzmann (MB) canonical distribution [49, 54], subject to energy conservation. We note that, for a uniform $\mathcal{P}_{0}(e)$, the distribution becomes independent of $s$. This situation corresponds to an infinite bath whose macroscopic state remains unchanged. Remarkably, in this formulation, the MB distribution emerges as a consequence of the FDR derived from the second law of thermodynamics, contrary to the conventional approach, where the FDR is obtained by assuming the MB distribution as the steady-state. Hence, the variational approach recovers the results of equilibrium statistical mechanics without being limited to it. The complete derivations supporting this section are presented in Appendix B.
III.4 Interconnected system of multiple reservoirs
@@
accordingly. The mass and entropy fluxes are given by $\mathcal{J}_{12}=-\mathcal{J}_{21}$ and $J_{12}=J_{21}$ respectively, with the stochastic counterparts given by the noises $\zeta(t),\eta(t)$ with amplitudes $\sigma=\sigma_{12}=-\sigma_{21}$ and $\kappa=\kappa_{12}=\kappa_{21}$. These symmetries derive from total mass and energy conservation due to the system being closed. For the purpose of the variational formulation, it is convenient to define the flux $J_{kl}$ for $k=l$ as [51]
$J_{kk}\coloneq-\sum_{l\neq k}J_{kl},$ (61)
@@
$\frac{\partial L}{\partial s_{k}}\delta\Sigma_{k}=\sum_{l}(J_{kl}+\kappa_{kl}\circ\eta)\delta\Gamma^{l}+\sum_{l}(\mathcal{J}_{l\rightarrow k}+\sigma_{l\rightarrow k}\circ\zeta)\delta W^{k}.$ (62)
@@
where the variables $\Gamma^{i},W^{i}$ are conjugate variables with $s_{i}$ (or $\Sigma_{i}$) and $N_{i}$, respectively. Taking variations then gives the following system:
@@
$\langle\bm{\xi}(t):\bm{\xi}(t^{\prime})\rangle=2\begin{bmatrix}1&C\\ C&1\end{bmatrix}\delta(t-t^{\prime}),$ (65)
where $C$ denotes the cross-correlation coefficient. It will be shown how cross-correlation implies a common hidden physical mechanism behind the heat and mass fluxes, giving rise to cross-effects (see Fig. 2).
Figure 2: Schematic of an interconnected system with two reservoirs exchanging heat ($J$) and mass ($\mathcal{J}$). Cross-effects induced by cross-correlated noise lead to additional fluxes $\mathcal{J}_{cross}$, enabling heat (mass) transfer in response to chemical potential (temperature) gradients.
@@
where the explicit expressions for $\nu_{m},\nu_{th}$ can be found in Appendix C. The average system and thermodynamic EPRs then yield, respectively,
$\dot{\mathcal{S}}=\int\frac{d\Omega}{(1-C^{2})}\left\{\frac{j^{2}_{N_{1}}}{\mathcal{P}\sigma^{2}}+\frac{j^{2}_{th}}{\mathcal{P}\kappa^{2}}+2C\frac{j_{N_{1}}j_{th}}{\mathcal{P}\sigma\kappa}+j_{N_{1}}\left(\frac{\tilde{\mathcal{J}}_{12}}{\sigma^{2}}-\frac{C}{\kappa\sigma}\tilde{J}_{12}\right)-j_{th}\left(\frac{\tilde{J}_{12}}{\kappa^{2}}-\frac{C}{\kappa\sigma}\tilde{\mathcal{J}}_{12}\right)\right\},$ (70)
$\sum_{k}\langle\!\langle\dot{s}_{k}\rangle\!\rangle=\int d\Omega\left\{j_{N_{1}}\left(\frac{\mu^{2}}{T^{2}}-\frac{\mu^{1}}{T^{1}}\right)+j_{th}\left(\frac{T^{1}-T^{2}}{T^{1}}+\frac{T^{2}-T^{1}}{T^{2}}\right)\right\}.$ (71)
Note that Eq. (52) is again satisfied since $\sum_{k}\langle\!\langle\dot{s}_{k}\rangle\!\rangle=\frac{d}{dt}\sum_{k}\langle s_{k}\rangle$ (see Appendix C). Adding both contributions and enforcing the second law gives the FDRs:
$\dot{S}_{tot}\geq 0\iff\begin{bmatrix}\vskip 4.2679pt\tilde{\mathcal{J}}_{12}\\ \tilde{J}_{12}\end{bmatrix}=\mathbb{L}\begin{bmatrix}\vskip 7.11317pt\displaystyle\frac{\mu^{1}}{T^{1}}-\frac{\mu^{2}}{T^{2}}\\ \displaystyle(T^{1}-T^{2})\left(\frac{1}{T^{1}}-\frac{1}{T^{2}}\right)\end{bmatrix}$, (72)
@@
being symmetric positive definite (PD) matrices. The average total EPR can then be written explicitly as a quadratic form:
$\dot{S}_{tot}=\int\frac{d\Omega}{\mathcal{P}}\begin{bmatrix}j_{N_{2}}&j_{th}\end{bmatrix}\mathbb{L}^{-1}\begin{bmatrix}j_{N_{2}}\\ j_{th}\end{bmatrix}\geq 0,$ (74)
with equality only at equilibrium, where dissipative currents vanish.
The FDR (72) shows that, due to cross-correlations, each flux is influenced by both temperature and chemical potential gradients, thereby exhibiting cross-effects. The relationship between the fluxes and the gradients is governed by the symmetric PD matrix $\mathbb{L}$, which embodies the Onsager reciprocal relations. This structure indicates that, in the macroscopic limit, the model inherently recovers the phenomenological relations of nonequilibrium thermodynamics [57, 51], mediated by the Onsager matrix. Namely, the FDR (72) has the same form as its macroscopic counterpart found in [51] from
@@
where $I$ denotes the macroscopic internal EP. In addition, the fluxes include noise-induced drifts, arising from multiplicative noise and/or state-dependent correlations, which are necessary for thermodynamic consistency. These features are particularly relevant in modeling active field theories, as discussed in Refs. [5, 6]. However, the special case $C=\pm 1$ leads to a degenerate situation in which the two noise terms become identical, introducing a redundancy and rendering the model ill-defined.
Since the system is closed, the functional
@@
satisfies the FP at steady state, analogously as in Sec. III.3, see Eq. (53). This can be shown by enforcing the FDRs (72) in the currents given by Eqs. (67), together with the relations
@@
Again, the system displays an equilibrium steady state under the FDRs derived through the second law, as we require for the thermodynamic consistency of closed systems. The functional (76) can be further simplified for specific forms of the energy $E(s_{1},s_{2},N_{1},N_{2})$ by applying the composite rule for Dirac delta functions given by Eq. (56).
For example, assuming that the reservoirs have equilibrated at a temperature $T=T^{1}=T^{2}$ and the grand canonical free energy $F(N_{1},N_{2})$ is separable, the energy function takes the form $E(s_{1},s_{2},N_{1},N_{2})=F(N_{1},N_{2})+T(s_{1}+s_{2})$, where $\partial_{N_{i}}F=\mu^{i}(N_{1},N_{2})$. The phase space integral $Z(e)$ then gives:
@@
Analogously to the example in Sec. III.3, this distribution favors states that minimize the total chemical free energy, subject to energy conservation. Assuming a uniform initial energy distribution, the prefactor can be absorbed into the normalization constant, giving the finite-dimensional analog distribution of Refs. [58, 59]. The complete derivations supporting this section are provided in Appendix C.
III.5 Mechanical open system
In this section, we present an example that illustrates the applicability of the framework to systems far from equilibrium. Specifically, we consider a particle driven by an external agent through a force $\mathbf{f}_{+}(\mathbf{x},\mathbf{p},\lambda)$, immersed in a heat bath undergoing external heat transfer. A key distinction from the previous examples is that the system is not constrained to relax to a steady state and may remain far from equilibrium throughout its evolution. The corresponding Lagrangian is given by

### v2

@@ III.4 Mechanical isolated system
where $\mathbf{f_{-}}=-\underline{\underline{\boldsymbol{\gamma}}}(\mathbf{x},\mathbf{p},s)\mathbf{p}/m$ is a dissipative force with state-dependent damping coefficient, $\underline{\underline{\mathbf{g}}}(\mathbf{x},\mathbf{p},s)$ is the noise strength and $\bm{\zeta}(t)$ is gaussian white noise verifying the relations given by Eq. (2). $\Gamma$ denotes the (generalized) thermal displacement, i.e. $\dot{\Gamma}=T$, an equality that follows from the variational principle.
@@
with $\dot{s}=\dot{\Sigma}$, equality that holds in any closed system. See Appendix C for the detailed derivation. Taking the total time derivative of $E=\frac{1}{2}m|\mathbf{v}|^{2}+U(\mathbf{x},s)$ we find the first law
@@
$-g^{2}\partial_{p}\mathcal{P}+\frac{pg^{2}}{mT}\partial_{s}\mathcal{P},$ (50a)
@@
the system and thermodynamic average EPRs yield (see Appendix C):
$\dot{\mathcal{S}}=\underbrace{\int d\Omega\,\bigg\{\frac{{j}_{d}^{2}}{\mathcal{P}g^{2}}}_{\langle\!\langle\dot{\mathfrak{s}}_{\rm tot}\rangle\!\rangle}$
@@
[…]
$\langle\!\langle\dot{\Sigma}\rangle\!\rangle=\langle\!\langle\dot{s}\rangle\!\rangle=-\int d\Omega\,\frac{p}{mT}{j}_{d}.$ (52)
The identification of the corresponding terms with $\langle\!\langle\dot{\mathfrak{s}}_{\rm tot}\rangle\!\rangle$ and $\langle\!\langle\dot{\mathfrak{s}}_{m}\rangle\!\rangle$ follows from a direct verification, see Appendix D. Adding both contributions gives the total EPR:
$\dot{S}_{\rm tot}$ $=\dot{\mathcal{S}}+\langle\!\langle\dot{\Sigma}\rangle\!\rangle=\int d\Omega\,\frac{{j}_{d}^{2}}{\mathcal{P}g^{2}}$ (53)
$-\int d\Omega\,{j}_{d}\left\{\frac{p}{mT}+\frac{1}{g^{2}}\left(-\gamma\frac{p}{m}-g\partial_{p}g+\frac{p}{m}g\partial_{s}(g/T)\right)\right\}.$
The first term is a PD quadratic form, but the second is sign-indefinite; hence, in contrast to standard ST, the total EPR need not be non-negative.
Enforcing the second law requires the vanishing of the term in braces multiplying $j_{d}$; see Appendix B. From this condition emerges the FDR and the associated total EPR:
$-\gamma\frac{p}{m}=-\frac{D}{T}\frac{p}{m}+g\partial_{p}g-g\partial_{s}(g/T)\frac{p}{m}\Rightarrow$ (54)
$\dot{S}_{\rm tot}=\int d\Omega\,\frac{{j}_{d}^{2}}{\mathcal{P}D}\geq 0,$ (55)
where we recall $D=g^{2}$. Notably, the total EPR has the same form as that obtained from the standard ST approach in one dimension, see Eq. (23), but it is instead defined over the thermodynamic phase space $\Omega$. The corresponding multidimensional version of Eq. (54), obtained by an analogous computation, is
@@
As anticipated in Sec. III.3, this solution, or equivalently, the associated FDR, enforces the identity $\dot{\Sigma}=\dot{\mathfrak{s}}_{m}$, see Eqs. (51) and (55). Hence, LDB [Eq. (45)] emerges directly from the second law axiom of the variational principle, without recourse to path probabilities. For completeness, Appendix D reproduces the same result via an explicit path probability calculation.
Remarkably, since the thermodynamic EP satisfies
@@
it implies that $\dot{S}_{\rm tot}$ is a total time derivative, as we can write
$\dot{S}_{\rm tot}=\frac{d}{dt}(\mathcal{S}+\langle s\rangle),$ (58)
in contrast with the standard ST setting, i.e. the results of Sec. II. Thus, for isolated systems, the entropy behaves as a function of state, as it is expected near equilibrium.
The FDR [Eq. (54)] connects the dissipative forces with the random forces covariance, plus contributions from noise-induced drift terms. In the special case of constant temperature and constant damping coefficient, we recover the standard Einstein relation discussed in Sec. II.3, namely $T\gamma=D$.
When $\gamma,T,g$ are independent of the entropy $s$, our result reproduces the FDR derived in Ref. [28], where this relation was derived by setting the steady-state distribution to be the Boltzmann distribution. This reference also illustrates how momentum-dependent multiplicative noise can arise, for instance, in systems with Coulomb friction, which requires an extension of the FDR to ensure thermodynamic consistency. Additionally, in this scenario—or more generally, for $\partial_{p}{\rm f}_{+}=0$—where the observable DoFs decouple from the entropy balance in (LABEL:eq:EoM_mech_closed), $\dot{s}$ recovers the same expression as $\dot{s}_{m}$ in Eq. (18) under the FDR constraint (56), providing it with a well-defined thermodynamic interpretation (Clausius relation).
Beyond recovering these known results, this framework introduces a novel approach for coupling mechanical and thermal DoFs [89]. This allows for taking into account finite baths and/or coupled system-bath dynamics. The third contribution in Eq. (54), the noise-induced drift due to $s$ dependence, is essential to guarantee thermodynamic consistency in the presence of such nonlinear system-bath coupling.
@@
which implies $\partial_{t}\mathcal{P}=-\nabla_{\Omega}\cdot\mathbf{j}=0$. The distribution (59) is general and holds for coupled system-bath dynamics as well. Moreover, it favors the states that maximize the entropy $s$, consistent with the maximum entropy principle in statistical mechanics, as well as being of the same form as the equilibrium distribution in fluctuation theory [14, 76, 56, 31], with an additional prefactor accounting for energy conservation.
@@
Due to the time-independence of $f(e)$, taking logarithms on both sides of Eq. (59) and taking the difference at $t_{1},t_{2}$ gives
$\Delta s=\Delta\ln\mathcal{P}=-\Delta s_{sys}\Rightarrow\Delta s_{\rm tot}=\Delta s_{sys}+\Delta s=0,$ (62)
where $s_{sys}$ is given by Eq. (35) and $s_{\rm tot}$ is the individual-trajectory level version of $S_{\rm tot}$. Thus the stochastic total entropy production vanishes along any trajectory at equilibrium.
@@
This corresponds precisely to the Maxwell-Boltzmann (MB) canonical distribution [14, 56], subject to energy conservation. We note that, for a uniform $\mathcal{P}_{0}(e)$, the distribution becomes independent of $s$. This situation corresponds to an infinite bath whose macroscopic state remains unchanged.
Remarkably, in this formulation, the MB distribution emerges as a consequence of the FDR derived from the second law of thermodynamics, contrary to the conventional approach, where the FDR is obtained by assuming the MB distribution as the steady-state. Hence, the variational approach recovers the results of equilibrium statistical mechanics in the canonical ensemble, without being restricted to such regime—recall we do not require the assumption of infinite baths for the FDRs nor the equilibrium ansatz [Eq. (59)].
III.5 Interconnected system of multiple reservoirs
@@
accordingly. The mass and entropy fluxes are given by $\mathcal{J}_{12}=-\mathcal{J}_{21}$ and $J_{12}=J_{21}$ respectively, with the stochastic counterparts given by the noises $\zeta(t),\eta(t)$ with amplitudes $\sigma=\sigma_{12}=-\sigma_{21}$ and $\kappa=\kappa_{12}=\kappa_{21}$. These symmetries are derived from total mass and energy conservation due to the system being closed. For the purpose of the variational formulation, it is convenient to define the flux $J_{kl}$ for $k=l$ as [42]
$J_{kk}:=-\sum_{l\neq k}J_{kl},$ (68)
@@
$\frac{\partial L}{\partial s_{k}}\delta\Sigma_{k}=\sum_{l}(J_{kl}+\kappa_{kl}\circ\eta)\delta\Gamma^{l}+\sum_{l}(\mathcal{J}_{lk}+\sigma_{lk}\circ\zeta)\delta W^{k}.$ (69)
@@
where the variables $\Gamma^{i},W^{i}$ are conjugate variables with $s_{i}$ (or $\Sigma_{i}$) and $N_{i}$, respectively. Taking variations then gives the following system (see Appendix E):
@@
$\langle\bm{\xi}(t)\otimes\bm{\xi}(t^{\prime})\rangle=2\begin{bmatrix}1&C\\ C&1\end{bmatrix}\delta(t-t^{\prime}),$ (72)
where $C$ denotes the cross-correlation coefficient. It will be shown how cross-correlation implies a common hidden physical mechanism behind the heat and mass fluxes, giving rise to cross-effects (see Fig. 2). The same reasoning applies to coupled-particle systems (see, e.g., [86, 6]), where analogous cross-effects arise naturally.
Figure 2: Schematic of an interconnected system with two reservoirs exchanging entropy ($J$) and mass ($\mathcal{J}$). Cross-effects induced by cross-correlated noise lead to additional fluxes $\mathcal{J}_{cross}$, enabling heat (mass) transfer in response to chemical potential (temperature) gradients.
@@
where the explicit expressions for $\nu_{m},\nu_{th}$ can be found in Appendix E. The average system and thermodynamic EPRs then yield, respectively,
$\dot{\mathcal{S}}=\int\frac{d\Omega}{\mathcal{P}}\begin{bmatrix}j_{N_{2}}\\ j_{th}\end{bmatrix}^{T}\mathbb{L}^{-1}\left\{\begin{bmatrix}j_{N_{2}}\\ j_{th}\end{bmatrix}-\begin{bmatrix}\tilde{\mathcal{J}}_{12}\mathcal{P}\\ \tilde{J}_{12}\mathcal{P}\end{bmatrix}\right\}=\langle\!\langle\dot{\mathfrak{s}}_{\rm tot}\rangle\!\rangle-\langle\!\langle\dot{\mathfrak{s}}_{m}\rangle\!\rangle,$ (77)
$\langle\!\langle\dot{\Sigma}\rangle\!\rangle=\sum_{k}\langle\!\langle\dot{s}_{k}\rangle\!\rangle=\int d\Omega\left\{j_{N_{2}}\left(\frac{\mu^{1}}{T^{1}}-\frac{\mu^{2}}{T^{2}}\right)+j_{th}\left(\frac{T^{1}-T^{2}}{T^{1}}+\frac{T^{2}-T^{1}}{T^{2}}\right)\right\},$ (78)
@@
being symmetric PD matrices. The identification of the corresponding terms with $\langle\!\langle\dot{\mathfrak{s}}_{\rm tot}\rangle\!\rangle$ and $\langle\!\langle\dot{\mathfrak{s}}_{m}\rangle\!\rangle$ follows from a direct verification, see Appendix E. Note that Eq. (58) is again satisfied since $\sum_{k}\langle\!\langle\dot{s}_{k}\rangle\!\rangle=\frac{d}{dt}\sum_{k}\langle s_{k}\rangle$, and thus $S_{\rm tot}$ behaves as a state function.
Adding both contributions, the average total EPR gives the form, with the shorthand notation $\bm{\mathfrak{J}}=(j_{N_{2}},j_{th})$:
$\dot{S}_{\rm tot}$ $=\dot{\mathcal{S}}+\langle\!\langle\dot{\Sigma}\rangle\!\rangle=\int\frac{d\Omega}{\mathcal{P}}\bm{\mathfrak{J}}^{T}\mathbb{L}^{-1}\bm{\mathfrak{J}}$
$-\int d\Omega\,\bm{\mathfrak{J}}^{T}\left\{\mathbb{L}^{-1}\begin{bmatrix}\tilde{\mathcal{J}}_{12}\\ \tilde{J}_{12}\end{bmatrix}-\begin{bmatrix}\vskip 5.69054pt\displaystyle\frac{\mu^{1}}{T^{1}}-\frac{\mu^{2}}{T^{2}}\\ \displaystyle\frac{T^{1}-T^{2}}{T^{1}}+\frac{T^{2}-T^{1}}{T^{2}}\end{bmatrix}\right\}$. (80)
Following the general argument in Appendix B, enforcing the second law requires that the term in braces vanish, yielding the FDRs and associated total EPR:
$\begin{bmatrix}\vskip 4.2679pt\tilde{\mathcal{J}}_{12}\\ \tilde{J}_{12}\end{bmatrix}=\mathbb{L}\begin{bmatrix}\vskip 7.11317pt\displaystyle\frac{\mu^{1}}{T^{1}}-\frac{\mu^{2}}{T^{2}}\\ \displaystyle(T^{1}-T^{2})\left(\frac{1}{T^{1}}-\frac{1}{T^{2}}\right)\end{bmatrix}\Rightarrow$ (81)
$\dot{S}_{\rm tot}=\int\frac{d\Omega}{\mathcal{P}}\bm{\mathfrak{J}}^{T}\mathbb{L}^{-1}\bm{\mathfrak{J}}\geq 0.$ (82)
The average total EPR is a PD quadratic form, attaining zero only at equilibrium where all dissipative currents vanish. Moreover, the FDRs enforce $\dot{\Sigma}=\dot{\mathfrak{s}}_{m}$, see Appendix E for an explicit path probability calculation. Hence, fully analogous to the mechanical example discussed in Sec. III.4, LDB [Eq. (45)] emerges directly from the second law axiom.
The matrix $\mathbb{L}$ has eigenvalues
which satisfy $\lambda_{\pm}>0$ for $|C|<1$, ensuring that $\mathbb{L}$ is PD. In this case, its inverse $\mathbb{L}^{-1}$ is also PD, since its eigenvalues are $1/\lambda_{\pm}>0$, which guarantees that $\dot{S}_{\rm tot}\geq 0$. The special case $C=\pm 1$ leads to a degenerate situation in which the two noise terms become identical, introducing a redundancy and rendering the model ill-defined.
The FDR (81) shows that, due to cross-correlations, each flux is influenced by both temperature and chemical potential gradients, thereby exhibiting cross-effects. The relationship between the fluxes and the gradients is governed by the symmetric PD matrix $\mathbb{L}$, which embodies the Onsager reciprocal relations. Indeed, the condition of LDB implies micro-reversibility, from which Onsager symmetry emerges [70, 33, 88, 74]. This structure indicates that, in the macroscopic limit, the model inherently recovers the phenomenological relations of nonequilibrium thermodynamics [70, 42], mediated by the Onsager matrix. Namely, the FDR (81) has the same form as its macroscopic counterpart found in [42] from
@@
where $I$ denotes the macroscopic internal EP. In addition, the fluxes include noise-induced drifts, arising from state-dependent correlations, which are necessary for thermodynamic consistency. These features are particularly relevant in modeling active field theories, as discussed in Refs. [12, 64]. Since the system is closed, the functional
@@
satisfies the FP at steady state, analogously as in Sec. III.4, see Eq. (59). This can be shown by enforcing the FDRs (81) in the currents given by Eqs. (74), together with the relations
@@
Again, the system displays an equilibrium steady state under the FDRs derived through the second law, as we require for the thermodynamic consistency of isolated systems. The functional (85) can be further simplified for specific forms of the energy $E(s_{1},s_{2},N_{1},N_{2})$ by applying the composite rule for Dirac delta functions given by Eq. (63).
For example, assuming that the reservoirs have equilibrated at a temperature $T=T^{1}=T^{2}$ and the Helmholtz free energy $F(N_{1},N_{2})$ is separable, the energy function takes the form $E(s_{1},s_{2},N_{1},N_{2})=F(N_{1},N_{2})+T(s_{1}+s_{2})$, where $\partial_{N_{i}}F=\mu^{i}(N_{1},N_{2})$. The phase space integral $Z(e)$ then gives:
@@
Analogously to the example in Sec. III.4, this distribution favors states that minimize the total chemical free energy, subject to energy conservation. Assuming a uniform initial energy distribution, the prefactor can be absorbed into the normalization constant, giving the finite-dimensional analog distribution of Refs. [25, 59].
III.6 Mechanical open system
In this section, we consider a particle driven by an external agent through a force $\mathbf{f}_{+}(\mathbf{x},\mathbf{p},\lambda)$, immersed in a heat bath undergoing external heat transfer; see, for instance, Refs. [90, 60, 22] for relevant experimental and numerical realizations. A key distinction from the previous examples is that the system is not constrained to relax to a steady state and may remain far from equilibrium throughout its evolution. The corresponding Lagrangian is given by